OpenStudy (anonymous):

Find the larger of two consecutive integers whose sum is 4 less than three times the smaller integer. how would I be to set up this problem.

OpenStudy (anonymous):

so: \[x + (x-1) = 3(x-1) - 4\]

OpenStudy (anonymous):

where x is the larger integer and x-1 is the smaller integer

OpenStudy (anonymous):

ok thanks ill tell you my anwser after i work it out will you tell me if i am correct?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is the second step 1x-1= 3x-3-4?

OpenStudy (anonymous):

the Left Hand Side is incorrect, look at your x's

OpenStudy (anonymous):

2x-1=3x-3-4

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

how would i get to the 3rd step because it is a werid looking problem to me haha

OpenStudy (anonymous):

first combine terms on the right hand side

OpenStudy (anonymous):

is it 2x-1=3x-7?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now get all of the x's on one side and all of the constants on the other

OpenStudy (anonymous):

5x=-6 or am i totally wrong?

OpenStudy (anonymous):

Be careful with your signs

OpenStudy (anonymous):

is is 2x-3x

OpenStudy (anonymous):

1x=-6?

OpenStudy (anonymous):

close!

OpenStudy (anonymous):

-1x=-6?

OpenStudy (anonymous):

yes, so x= ?

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

\[-x = -6\] divide by -1 on both sides: \[x = 6\]

OpenStudy (anonymous):

when ever you divide by to negitives you get a positive i forgot that thanks :) so do i do 6-1for the seccond number ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it would be 5,6?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks you so much i really apricaite you taking your time to help me :)

OpenStudy (anonymous):

You are welcome